# Lidan Shou1,2, Gang Chen1,2, Huan Li1,2✉ 1

### v1

@@ Abstract
Yicheng Ji1,2✌, Jun Zhang1,2✌, Heming Xia3, Jinpeng Chen4, Lidan Shou1,2, Gang Chen1,2, Huan Li1,2✉ 1The State Key Laboratory of Blockchain and Data Security, Zhejiang University 2Hangzhou High-Tech Zone (Binjiang) Institute of Blockchain and Data Security 3Department of Computing, The Hong Kong Polytechnic University 4School of Computer Science, Beijing University of Posts and Telecommunications {jiyicheng.cs, zj.cs, should, cg, lihuan.cs}@zju.edu.cn, heming.xia@connect.polyu.hk, jpchen@bupt.edu.cn
@@ 1 Introduction
Video large language models (Vid-LLMs) (Li et al., 2024a; Ahmed et al., 2025; Lin et al., 2024; Zhou et al., 2025) have demonstrated strong performance in video comprehension. Most Vid-LLMs use a sequential visual representation, encoding sampled frames into tens of thousands of video tokens alongside language prompts to achieve high generation performance. However, as long videos become more common, this design introduces significant memory and computational costs. For instance, LLaVA-OneVision (Li et al., 2024a) processes each video frame into 196 tokens, meaning a two-minute video at 60 FPS would require more than 1 million tokens by default without any reduction. The large number of video tokens increases the sequence length, resulting in quadratic attention overhead during prefilling. During decoding, the autoregressive nature of generation exacerbates the memory-bound issue, as the growing key-value (KV) cache must be loaded and stored in GPU memory alongside model parameters, limiting scalability and increasing latency (Lin et al., 2024).
@@
Fortunately, speculative decoding (SD) offers a promising solution to accelerate LLM decoding without sacrificing quality (Leviathan et al., 2023) by using a lightweight draft model to propose multiple draft tokens, which are then verified in parallel by a target model. However, deploying SD for Vid-LLMs is challenging. First, autoregressive draft models for LLMs (Du et al., 2024; Li et al., 2024d, c) suffer from reduced efficiency in video scenarios because they require linearly increasing KV caches, which become the dominant bottleneck as the length of the video grows. Second, the visual context in video is relatively long and sparse, with significant redundancy. While existing SD methodologies tailored for long-context scenarios (Sun et al., 2024; Chen et al., 2025; Yang et al., 2025a) are modality-unaware, they fail to exploit the heavy redundancy and distinct attention patterns of video tokens (detailed in Appendix˜C), leading to performance degradation (see Section˜4.3). These gaps motivate us to perform video token pruning for the draft model, thereby reducing its KV cache size and enhancing speculation efficiency.
@@ 3.2 Two-Stage Token Pruning for Draft Model
Stage I: Top-P Retention of Highly Informative Tokens
@@ 4 Experiments
Since SpecVLM mainly focuses on the acceleration during decoding, we evaluate its performance in video captioning and video description tasks, which require generating long text paragraphs. The benchmarks we use include VideoDetailCaption (LMMs-Lab, 2024), MVBench (Li et al., 2024b), MVLU (Zhou et al., 2025), and LongVideoBench (Wu et al., 2024). Experimental details are elaborated in Appendix˜A.
@@ 4.4 Exploration of Diverse Pruning Criteria
4.5 Impact of Decoding Steps
4.6 Breakdown of Computation
Apart from accelerating the decoding stage of target Vid-LLMs, SpecVLM introduces minimal overhead in the pruning process, as illustrated in Table˜4. Owing to video token pruning, the prefill length of the draft model is substantially reduced, and the additional prefilling time becomes negligible compared to the target model inference time. In this work, we significantly reduce the target model’s decoding time—a level of efficiency that cannot be achieved by prior methods relying solely on token reduction.
5 Related Work
5.1 Speculative Decoding
Speculative Decoding for LLMs.
Speculative decoding is shown to be an effective approach to accelerate LLMs while maintaining the original output distribution. It relies on two key processes: efficient drafting and parallel verification. Initial explorations (Leviathan et al., 2023; Guo et al., 2023; Kim et al., 2023; Xia et al., 2023) attempt to use existing small LLMs as draft models to ensure reliable speculation. Self-speculative methods (Xia et al., 2025; Zhang et al., 2024a) use partial layers of the original model to generate predictions, without introducing extra models. Recent advancements (Cai et al., 2024; Li et al., 2024d; Du et al., 2024) focus on enhancing the efficiency of the drafting stage. These works attain high acceleration ratios through a specially trained draft model with reduced latency. Meanwhile, tree-based speculation methods (Li et al., 2024c, d; Miao et al., 2024) are proposed to boost the average accept length, by predicting multiple candidates and forming draft trees. Among the aforementioned draft models, the success of the previous state-of-the-art EAGLE method (Li et al., 2024c) highlights the autoregressive structure as a key factor in improving the accuracy of the speculation. However, autoregressive draft models need to maintain their own KV caches, which introduces additional memory overhead when faced with long video input.
Long-Context Speculative Decoding.
Long-context speculative decoding methodologies (Sun et al., 2024; Chen et al., 2025; Yang et al., 2025a) are specially developed to mitigate the substantial overhead of KV cache. Specifically, TriForce (Sun et al., 2024) introduces a hierarchical speculation to tackle the two bottlenecks of model weights and KV cache. MagicDec (Chen et al., 2025) reassesses the trade-off between throughput and latency. LongSpec (Yang et al., 2025a) proposes a memory-efficient draft model with a constant memory footprint. Appendix˜C elaborates on the reasons why these methods are not suitable for Vid-LLMs.
5.2 Visual Token Reduction
Compared to information-dense text, visual tokens often exhibit high redundancy, making token reduction an effective way to reduce computational and memory overhead. Recent studies largely focus on training-free visual token reduction methods based on token importance and redundancy. FastV (Chen et al., 2024a) selects important visual tokens after layer 2 using attention scores of MLLMs. While SparseVLM (Zhang et al., 2024b) evaluates the visual relevance of text tokens and performs pruning based on the attention scores of a subset of text tokens. VisionZip (Yang et al., 2025b) reduces visual redundancy in the vision encoders. DART (Wen et al., 2025) prunes tokens based on its duplication with other tokens. Currently, video token reduction draws increasing attention due to the high volume of video tokens. DyCoke (Tao et al., 2025) performs token merging across frames and reduces KV cache dynamically. PrunVID (Huang et al., 2025) identifies static and dynamic tokens across frames, and selectively prunes visual features relevant to question tokens. FastVID (Shen et al., 2025) partitions frames into segments and introduces a density-based token pruning strategy. Notably, recent Tang et al. (2025) applies temporal processing during video sampling by performing key frame extraction, whereas ours prunes visual tokens post-encoder, making the two methods orthogonal.
6 Conclusion
[…]
7 Ethical Considerations
8 Limitation
While our enhanced speculative decoding framework brings clear benefits for accelerating Vid-LLMs, there are a few limitations to consider. First, we introduce an additional draft model during inference. Although its computational overhead is relatively small compared to the target model (or can be avoided using the Self-SD setting), the choice of the draft model requires careful consideration to achieve optimal speedup. Second, to avoid introducing training as an additional variable, we use an existing Vid-LLM as the draft model without further fine-tuning. This design choice imposes certain constraints on the maximum achievable acceleration. Nevertheless, our method has the potential to be seamlessly integrated with smaller and faster draft models, as it only requires a one-time pruning of the draft model’s KV cache during the prefilling stage. As the technology for training lightweight Vid-LLMs is still in its early stages and lacks suitable candidates, we believe future work can explore more draft model optimization techniques (Li et al., 2024d; Du et al., 2024) to address this limitation.
Appendix A Experimental Details
Task and Benchamark Details.
Since SpecVLM mainly focuses on the lossless acceleration during the decoding stage, we select video captioning and video description as our tasks, which require the model to summarize and describe the video by understanding detailed subjects and actions, and involve generating relatively long text paragraphs. Unlike conventional video question answering (VQA), we prompt the model to generate a description rich in visual information instead of direct answer selection. We sample videos from mainstream video understanding benchmarks, including VideoDetailCaption (LMMs-Lab, 2024), MVBench (Li et al., 2024b), MVLU (Zhou et al., 2025), and LongVideoBench (Wu et al., 2024), to ensure a comprehensive coverage of different durations and varying scenarios. For each benchmark, 50 instances are randomly sampled. For LLaVA-OneVision-72B and LLaVA-OneVision-7B, we uniformly sample 64 and 128 frames to generate a 196 $\times$ 64 and 196 $\times$ 128 video token input, respectively. For Qwen2.5-VL series, we adjust the FPS to generate input of comparable length.
Implementation Details.
Appendix B Impact of Tree Attention
As illustrated in Table˜5, using a tree structure for drafting and verification greatly enhances the average accept length by validating multiple candidates in a single forward. When incorporated with tree attention, SpecVLM achieves a higher speculation accuracy. The draft tree structure adopted in SpecVLM is depicted as Fig.˜9. The chosen tree structure is motivated by the intuition that candidate tokens with higher probabilities merit deeper and wider expansion, whereas low-probability tokens should not be further explored.
Method $\tau$
SD-Chain 3.12
SD-Tree 3.68
SpecVLM-Chain 2.79
SpecVLM 3.48
@@ Appendix B Impact of Tree Attention
Appendix C Why is Long-context SD Technologies Not Suitable for Vid-LLMs?
@@
Appendix D Discussion on Smaller Draft Model
In this section, we discuss the effectiveness of SpecVLM in a smaller draft model scenario. Although the minimum draft model used in our experiment is 7b, we believe our method still has benefits for smaller draft models from the following two perspectives. (i) SpecVLM prunes redundant video tokens at inference time, reducing the training cost and architectural complexity of high-quality small draft models for video understanding. Existing small draft models for LLMs like EAGLE do not support long-context input sequences. Training small draft models to match the context length of Vid-LLMs would require massive computation, and it remains questionable whether these small models can achieve the same accuracy at video input context lengths (Sun et al., 2024). Therefore, reducing the input length of the small draft model through our video token pruning strategy allows it to perform efficient speculation without the need for costly long-context training. (ii) For smaller draft models, the overhead caused by input length still persists. To simulate a scenario similar to EAGLE, we tested the latency of a single layer of the LLaVA-OneVision-0.5B model at varying input lengths. Latency is calculated on average of 100 decoded tokens on a single Nvidia A100 GPU. The results in Table˜6 show that the latency of the small draft model increases significantly with input length, eventually becoming multiple times larger than the original. Given the increasing trend of video inputs (e.g., Long video input with million-level video tokens (Chen et al., 2024b) ), the additional draft overhead from the KV Cache would gradually undermine the lightweight nature of smaller draft models. In other words, combining draft models with token pruning still provides a performance benefit.
@@ Appendix D Discussion on Smaller Draft Model
Appendix E Case Study

### v2

# Lidan Shou1,2, Gang Chen1, Huan Li1,2✉ 1

@@ Abstract
Yicheng Ji1,2✌, Jun Zhang1,2✌, Heming Xia3, Jinpeng Chen4, Lidan Shou1,2, Gang Chen1, Huan Li1,2✉ 1The State Key Laboratory of Blockchain and Data Security, Zhejiang University 2Hangzhou High-Tech Zone (Binjiang) Institute of Blockchain and Data Security 3Department of Computing, The Hong Kong Polytechnic University 4School of Computer Science, Beijing University of Posts and Telecommunications {jiyicheng.cs, zj.cs, should, cg, lihuan.cs}@zju.edu.cn, heming.xia@connect.polyu.hk, jpchen@bupt.edu.cn
@@ 1 Introduction
Video large language models (Vid-LLMs) (Li et al., 2024a; Ahmed et al., 2025; Lin et al., 2024; Fang et al., 2024) have demonstrated strong performance in video comprehension. Most Vid-LLMs use a sequential visual representation, encoding sampled frames into tens of thousands of video tokens alongside language prompts to achieve high generation performance. However, as long videos become more common, this design introduces significant memory and computational costs. For instance, LLaVA-OneVision (Li et al., 2024a) processes each video frame into 196 tokens, meaning a two-minute video at 60 FPS would require more than 1 million tokens by default without any reduction. The large number of video tokens increases the sequence length, resulting in quadratic attention overhead during prefilling. During decoding, the autoregressive nature of generation exacerbates the memory-bound issue, as the growing key-value (KV) cache must be loaded and stored in GPU memory alongside model parameters, limiting scalability and increasing latency (Lin et al., 2024).
@@
Fortunately, speculative decoding (SD) offers a promising solution to accelerate LLM decoding without sacrificing quality (Leviathan et al., 2023) by using a lightweight draft model to propose multiple draft tokens, which are then verified in parallel by a target model. However, deploying SD for Vid-LLMs is challenging. First, autoregressive draft models for LLMs (Du et al., 2024; Li et al., 2024d, c) suffer from reduced efficiency in video scenarios because they require linearly increasing KV caches, which become the dominant bottleneck as the length of the video grows. Second, the visual context in video is relatively long and sparse, with significant redundancy. While existing SD methodologies tailored for long-context scenarios (Sun et al., 2024; Chen et al., 2025; Yang et al., 2025a) are modality-unaware, they fail to exploit the heavy redundancy and distinct attention patterns of video tokens (detailed in Appendix˜D), leading to performance degradation (see Section˜4.3). These gaps motivate us to perform video token pruning for the draft model, thereby reducing its KV cache size and enhancing speculation efficiency.
@@ 3.2 Two-Stage Token Pruning for Draft Model
Stage I: Top-P Retention of Highly Informative Tokens.
@@ 4 Experiments
Since SpecVLM mainly focuses on the acceleration during decoding, we evaluate its performance in video captioning and video description tasks, which require generating long text paragraphs. The benchmarks we use include VideoDetailCaption (LMMs-Lab, 2024), MVBench (Li et al., 2024b), MVLU (Zhou et al., 2025a), and LongVideoBench (Wu et al., 2024). Experimental details are elaborated in Appendix˜A.
@@ 4.4 Exploration of Diverse Pruning Criteria
4.5 Impact of Decoding Steps
5 Related Work
5.1 Speculative Decoding
Speculative Decoding for LLMs.
Speculative decoding is shown to be an effective approach to accelerate LLMs while maintaining the original output distribution. It relies on two key processes: efficient drafting and parallel verification. Initial explorations (Leviathan et al., 2023; Guo et al., 2023; Kim et al., 2023; Xia et al., 2023) attempt to use existing small LLMs as draft models to ensure reliable speculation. Self-speculative methods (Xia et al., 2025; Zhang et al., 2024a; Song et al., 2025) use partial layers of the original model to generate predictions, without introducing extra models. Recent advancements (Cai et al., 2024; Li et al., 2024d; Du et al., 2024) focus on enhancing the efficiency of the drafting stage. These works attain high acceleration ratios through a specially trained draft model with reduced latency. Meanwhile, tree-based speculation methods (Li et al., 2024c, d; Miao et al., 2024) are proposed to boost the average accept length, by predicting multiple candidates and forming draft trees. Among the aforementioned draft models, the success of the previous state-of-the-art EAGLE method (Li et al., 2024c) highlights the autoregressive structure as a key factor in improving the accuracy of the speculation. However, autoregressive draft models need to maintain their own KV caches, which introduces additional memory overhead when faced with long video input.
Long-Context Speculative Decoding.
Long-context speculative decoding methodologies (Sun et al., 2024; Chen et al., 2025; Yang et al., 2025a) are specially developed to mitigate the substantial overhead of KV cache. Specifically, TriForce (Sun et al., 2024) introduces a hierarchical speculation to tackle the two bottlenecks of model weights and KV cache. MagicDec (Chen et al., 2025) reassesses the trade-off between throughput and latency. LongSpec (Yang et al., 2025a) proposes a memory-efficient draft model with a constant memory footprint. Appendix˜D elaborates on the reasons why these methods are not suitable for Vid-LLMs.
5.2 Visual Token Reduction
Compared to information-dense text, visual tokens often exhibit high redundancy, making token reduction an effective way to reduce computational and memory overhead. Recent studies largely focus on training-free visual token reduction methods based on token importance and redundancy. FastV (Chen et al., 2024a) selects important visual tokens after layer 2 using attention scores of MLLMs. While SparseVLM (Zhang et al., 2024b) evaluates the visual relevance of text tokens and performs pruning based on the attention scores of a subset of text tokens. VisionZip (Yang et al., 2025b) reduces visual redundancy in the vision encoders. DART (Wen et al., 2025) prunes tokens based on its duplication with other tokens. Currently, video token reduction draws increasing attention due to the high volume of video tokens. DyCoke (Tao et al., 2025) performs token merging across frames and reduces KV cache dynamically. PrunVID (Huang et al., 2025) identifies static and dynamic tokens across frames, and selectively prunes visual features relevant to question tokens. FastVID (Shen et al., 2025) partitions frames into segments and introduces a density-based token pruning strategy. VidCom2 (Liu et al., 2025) dynamically adjusts compression intensity based on frame uniqueness. Notably, recent Tang et al. (2025) applies temporal processing during video sampling by performing key frame extraction, whereas ours prunes visual tokens post-encoder, making the two methods orthogonal.
6 Conclusion
[…]
7 Acknowledgements
The work was supported by the Major Research Program of Zhejiang Provincial Natural Science Foundation (No. LD24F020015), Zhejiang Province "Leading Talent of Technological Innovation Program" (No. 2023R5214), Beijing Natural Science Foundation(Grant No.L233034), and Fundamental Research Funds for the Beijing University of Posts and Telecommunications (Grant No.2025TSQY01).
8 Ethical Considerations
9 Limitation
While our enhanced speculative decoding framework brings clear benefits for accelerating Vid-LLMs, there are a few limitations to consider. First, our method is primarily applicable to resource-constrained long-video scenarios, where memory bandwidth becomes the dominant bottleneck. Second, we introduce an additional draft model during inference. Although its computational overhead is relatively small compared to the target model (or can be avoided using the Self-SD setting), the choice of the draft model requires careful consideration to achieve optimal speedup. Moreover, to avoid introducing training as an additional variable, we use an existing Vid-LLM as the draft model without further fine-tuning. This design choice imposes certain constraints on the maximum achievable acceleration. Nevertheless, our method has the potential to be seamlessly integrated with smaller and faster draft models, as it only requires a one-time pruning of the draft model’s KV cache during the prefilling stage. As the technology for training lightweight Vid-LLMs is still in its early stages and lacks suitable candidates, we believe future work can explore more draft model optimization techniques (Li et al., 2024d; Du et al., 2024) to address this limitation.
Appendix A Experimental Details
Task and Benchamark Details.
Since SpecVLM mainly focuses on the lossless acceleration during the decoding stage, we select video captioning and video description as our tasks, which require the model to summarize and describe the video by understanding detailed subjects and actions, and involve generating relatively long text paragraphs. Unlike conventional video question answering (VQA), we prompt the model to generate a description rich in visual information instead of direct answer selection. We sample videos from mainstream video understanding benchmarks, including VideoDetailCaption (LMMs-Lab, 2024), MVBench (Li et al., 2024b), MVLU (Zhou et al., 2025a), and LongVideoBench (Wu et al., 2024), to ensure a comprehensive coverage of different durations and varying scenarios. For each benchmark, 50 instances are randomly sampled. For LLaVA-OneVision-72B and LLaVA-OneVision-7B, we uniformly sample 64 and 128 frames to generate a 196 $\times$ 64 and 196 $\times$ 128 video token input, respectively. For Qwen2.5-VL series, we adjust the FPS to generate input of comparable length.
Implementation Details.
Appendix B Impact of Tree Attention
As illustrated in Table˜4, using a tree structure for drafting and verification greatly enhances the average accept length by validating multiple candidates in a single forward. When incorporated with tree attention, SpecVLM achieves a higher speculation accuracy. The draft tree structure adopted in SpecVLM is depicted as Fig.˜9. The chosen tree structure is motivated by the intuition that candidate tokens with higher probabilities merit deeper and wider expansion, whereas low-probability tokens should not be further explored.
Method $\tau$
SD-Chain 3.12
SD-Tree 3.68
SpecVLM-Chain 2.79
SpecVLM 3.48
@@ Appendix B Impact of Tree Attention
Appendix C Breakdown of Computation
Apart from accelerating the decoding stage of target Vid-LLMs, SpecVLM introduces minimal overhead in the pruning process, as illustrated in Table˜5. Owing to video token pruning, the prefill length of the draft model is substantially reduced, and the additional prefilling time becomes negligible compared to target model inference time. In this work, we significantly reduce the target model’s decoding time—a level of efficiency that cannot be achieved by prior methods relying solely on token reduction.
Appendix D Why is Long-context SD Technologies Not Suitable for Vid-LLMs?
@@
Appendix E Discussion on Smaller Draft Model
In this section, we discuss the effectiveness of SpecVLM in a smaller draft model scenario. Although the minimum draft model used in our experiment is 7b, we believe our method still has benefits for smaller draft models from the following two perspectives. (i) SpecVLM prunes redundant video tokens at inference time, reducing the training cost and architectural complexity of high-quality small draft models for video understanding. Existing small draft models for LLMs like EAGLE do not support long-context input sequences. Training small draft models to match the context length of Vid-LLMs would require massive computation, and it remains questionable whether these small models can achieve the same accuracy at video input context lengths (Sun et al., 2024). Therefore, reducing the input length of the small draft model through our video token pruning strategy allows it to perform efficient speculation without the need for costly long-context training. (ii) For smaller draft models, the overhead caused by input length still persists. To simulate a scenario similar to EAGLE, we tested the latency of a single layer of the LLaVA-OneVision-0.5B model at varying input lengths. Latency is calculated on average of 100 decoded tokens on a single Nvidia A100 GPU. The results in Table˜6 show that the latency of the small draft model increases significantly with input length, eventually becoming multiple times larger than the original. Given the increasing trend of video inputs (e.g., Long video input with million-level video tokens (Chen et al., 2024b) ), the additional draft overhead from the KV Cache would gradually undermine the lightweight nature of smaller draft models. In other words, combining draft models with token pruning still provides a performance benefit. In future work, we will investigate memory- and data-efficient training strategies for Vid-LLMs (Li et al., 2024d; Hu et al., 2022; Zhang et al., 2025b, c), with the goal of constructing more optimized draft models that remain effective in memory-constrained devices Zhou et al. (2025b) and scalable to multi-tenant settings Zhang et al. (2025a).
@@ Appendix E Discussion on Smaller Draft Model
Appendix F Case Study